IT took 3 hours to row a boat 18 km against the current. The return trip with the current took1 1/2 hours. Find the speed of the rowboat in still water.
Answer by Alan3354(69443)

IT took 3 hours to row a boat 18 km against the current. The return trip with the current took1 1/2 hours. Find the speed of the rowboat in still water.
-------------
18/3 = 6 km/hr
18/1.5 = 12 km/hr
--------------
The boat's speed is the avg = (6 + 12)/2
= 9 km/hr